The charges for a long distance telephone call is $125 for the first five minutes and $32 for each additional minutes or fraction of a minute. What is the largest whole number of minutes a call could last if it cost you less than $1200?
Answer by ankor@dixie-net.com(22740)

The charges for a long distance telephone call is $125 for the first five minutes and $32 for each additional minutes or fraction of a minute.
What is the largest whole number of minutes a call could last if it cost you less than $1200?
let m = total no. of minutes used
32(m-5) + 125 = 1200
32m - 160 = 1200 - 125
32m = 1075 + 160
32m = 1235
m = 1235/32
m = 38 whole min
:
Check the cost using 38 min of use
(32(38-5) + 125 = $1181